OpenStudy (anonymous):

The inequality x + 2y ≥ 3 is satisfied by point (1, 1). True False

OpenStudy (anonymous):

Plug in and try it out :-)

OpenStudy (anonymous):

it is false

OpenStudy (callisto):

@sexy1993 It's true... and you're wrong! Consider LHS x+ 2y = 1+2(1) = 3, which satisfy the inequality...

OpenStudy (anonymous):

(1,1) is true. in case { x+2y = 3} and false in case { x+2y >3}